A point source emits light of only a single frequency = 6.7 x 1014 Hz. What color is the light produced?
1
Expert's answer
2020-12-03T06:54:35-0500

It is easier to figure out what color it is by the wavelength, so find it:


"\lambda=\frac{c}{f}=\frac{3\cdot10^8}{6.7\cdot10^{14}}=448\text{ nm}."

448 nm is the wavelength of blue color.